OpenStudy (anonymous):

Help. I'm learning about radical equations and I;m confused about how √U = -8 is undefined. Can someone explain to me why please.

OpenStudy (anonymous):

@Nnesha

OpenStudy (anonymous):

hello

Nnesha (nnesha):

ello o^_^o

Nnesha (nnesha):

os is it \[\huge\rm \sqrt{U}= \sqrt{-8} \] ?

OpenStudy (anonymous):

wait what?

Nnesha (nnesha):

sqrt u = sqrt -8 ?

OpenStudy (anonymous):

I don't think so

OpenStudy (anonymous):

I thought you square both sides to get rid of the square symbol

OpenStudy (anonymous):

and it gets me \[U = 64\]

Nnesha (nnesha):

no i'm asking a question so question is \[\sqrt{u}= -8\]

OpenStudy (anonymous):

root symbol*

OpenStudy (anonymous):

yes

Nnesha (nnesha):

okay yes take square both side to get rid of square root

OpenStudy (anonymous):

and it got me \[U = 8\] but the book says it has no solution

OpenStudy (anonymous):

whoops

OpenStudy (anonymous):

\[U=64\]

Nnesha (nnesha):

yes right take square both side \[(\sqrt{u})^2 = (-8)^2\]

Nnesha (nnesha):

yep right now 64 is an extraneous solution meaning when you plug it back you wouldn't get equal sides

Nnesha (nnesha):

replace u by 64 in the original equation

OpenStudy (anonymous):

Oh I think I get it now

Nnesha (nnesha):

yep \[\huge\rm \sqrt{64}= -8 \] ? solve it both sides are equal ?

OpenStudy (anonymous):

Nope

OpenStudy (anonymous):

\[8 = -8\]

OpenStudy (anonymous):

Ok, so thats why it says its no solution than

Nnesha (nnesha):

so yes both sides are not equal that's why no solution 64 is an extraneous solution

OpenStudy (anonymous):

Thank you so much! That really helped me out.

Nnesha (nnesha):

my pleasure :-)